Circular No. 7504

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dq IN MCG +00-6-43 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova (mag ∼ 18.5) on unfiltered images taken on Oct. 1.4 and 3.3 UT. An earlier image taken on Sept. 19.4 also showed a hint of the new object. SN 2000dq is located at α = 2^{h}08^{m}20^{s}.98, δ = +1^{o}11′00′′.8 (equinox 2000.0), which is 11′′.4 east and 2′′.2 south of the nucleus of MCG +00-6-43. A KAIT image of the same field on Sept. 12.4 showed nothing at this position (limiting mag ∼ 19.0).

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000dq, obtained by P. Berlind on Oct. 4.36 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova in the first week after maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen and Ca II. The absorption of the Hα line has a double minimum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 12 687 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9200 km/s.

2000 DP_{107}

P. Pravec and P. Kušnirák, Ondřejov Observatory; M. Hicks, Table Mountain Observatory; B. Holliday, Corpus Christi Observatory; and B. Warner, Palmer Divide Observatory, report: “Our photometric observations of this object (cf. IAUC 7496, 7503) on nine nights during Sept. 25.2–Oct. 4.2 UT revealed a lightcurve that shows a fast variation with a period of 0.11564 ± 0.00001 day and an amplitude of 0.19 mag, which we intepret as being a rotation lightcurve of the primary, and there occurred brightness attenuations with a period of 1.76 ± 0.02 days that we interpret as eclipse/occultation events in the binary system. The estimated secondary-to-primary diameter ratio is 0.37 (uncertainty ∼ 10 percent). The measured color indices, V − R = +0.388 ± 0.013, B − R = +1.088 ± 0.019, U − R = +1.366 ± 0.025, and R − I = +0.329 ± 0.011, reveal that 2000 DP_{107} is spectrally flat, consistent with a C-type minor planet. The next eclipse/occultation events are predicted to occur around Oct. 5.18, 6.04, 6.94, and 7.80 UT. Details and updates are available at

http://sunkl.asu.cas.cz/~ppravec/2000DP107.htm.”

2000 October 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green